OpenStudy (anonymous):

@TheSmartOne find the maximum profit if P = 4x + 5y.

OpenStudy (danjs):

is there more to this question?

OpenStudy (danjs):

is this a part of a larger question or something?

OpenStudy (anonymous):

Kind of, I don't see how it relates though. Graph the feasibility region represented by the given inequalities and find the maximum profit if P = 4x + 5y. x ≥ 5 y ≤ 10 2x - y ≤ 8

OpenStudy (anonymous):

\[x \ge 5\]

OpenStudy (anonymous):

\[y \le10\]

OpenStudy (anonymous):

\[2x-y \le8\]

OpenStudy (danjs):

The process is: a) Plot the graph for each b) Find the four corner points c) Find the profit for each Point d) What is the maximum profit?

OpenStudy (danjs):

Here is the graph of this problem

OpenStudy (danjs):

find the profit for those points, which is the max

OpenStudy (danjs):

Profit = 4x + 5y Try points (x,y) = (5,10), (5,2), and (9,10) ...which has the highest profit

OpenStudy (anonymous):

(9, 10)?

OpenStudy (danjs):

U understand? the profit is restricted by those 3 given conditions, x is greater or equal to 5, y is less or equal to 10, and 2x-y is less than or equal to 8

OpenStudy (danjs):

the profit has to be contained in the overlapping region of those 3 inequalities, the triangle formed on that linked graph..

OpenStudy (danjs):

Try, (x,y) = (9,10) Profit = 4*9 + 5*10 = 36 + 50 = 86

OpenStudy (danjs):

yep, that is greater than the other points

OpenStudy (anonymous):

Nevermind, I misread. I think your answer's correct. Thank you.

OpenStudy (danjs):

no prob, i think the main point of the prob, is to realize the restriction on the x and y values for the profit function... Which is the triangle formed by the overlap of the 3 graphs.

OpenStudy (danjs):

and the maximum and minimum values, will be somewhere on those vertices of the triangle

OpenStudy (anonymous):

The graph that I got and you got a different.

OpenStudy (anonymous):

Nevermind, I took the wrong part of the graph.

OpenStudy (danjs):

That is the same graph, make the view window larger

OpenStudy (danjs):

If ya get stuck in the future, you can tag me to a question with @DanJS if you want